# Supersonic–sonic patch solution for the two-dimensional Euler equations with a van der Waals equation of state 11footnotemark: 1

### v1

@@ 3.5 Regularity of the solution
$\underline{K}t_{m}\leq|z_{2}-z_{1}|^{1/3}\leq\overline{K}t_{m},$ (3.46)
for some uniform positive constants $\underline{K}$ and $\overline{K}$. Since $\widetilde{W}$ is uniformly bounded, we may then apply (3.23) together with Lemma 3.3 to obtain
$|\tilde{\partial}_{+}\widetilde{X}|,|\tilde{\partial}_{-}\widetilde{Y}|,|\widetilde{X}_{t}|,|\widetilde{Y}_{t}|,t^{1/2}|\widetilde{X}_{z}|,t^{1/2}|\widetilde{Y}_{z}|\leq\overline{M_{1}},$ (3.47)
for some positive constant $\overline{M}_{1}>0$. Therefore, by Corollary 3.2 together with (3.46) and (3.47), we obtain
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|&=|\widetilde{Y}(z_{1},0)-\widetilde{X}(z_{2},0)|\\ &\leq|\widetilde{Y}(z_{1},0)-\widetilde{Y}(z_{m},t_{m})|+|\widetilde{Y}(z_{m},t_{m})-\widetilde{X}(z_{m},t_{m})|+|\widetilde{X}(z_{m},t_{m})-\widetilde{X}(z_{2},0)|\\ &\leq\overline{M_{1}}t_{m}+2\sup_{\Omega}|\widetilde{W}|\cdot t_{m}+\overline{M_{1}}t_{m}\\ &\leq\frac{2}{\overline{K}}(\overline{M_{1}}+\sup_{\Omega}|\widetilde{W}|)|z_{2}-z_{1}|^{1/3}.\end{split}$ (3.48)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{2})|&=|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{1})|+|\widetilde{X}(z_{2},t_{1})-\widetilde{X}(z_{2},t_{2})|\\ &\leq\sup_{\Omega}|\widetilde{X}_{z}|\cdot|z_{1}-z_{2}|+\sup_{\Omega}|\widetilde{X}_{t}|\cdot|t_{1}-t_{2}|\\ &\leq\frac{\overline{M_{1}}}{t^{1/2}}|z_{1}-z_{2}|+\overline{M_{1}}|t_{1}-t_{2}|\\ &\leq\overline{M_{1}}|z_{1}-z_{2}|^{1/2}+\overline{M_{1}}|t_{1}-t_{2}|\\ &\leq 2\overline{M_{1}}|(z_{1},t_{1})-(z_{2},t_{2})|^{1/3},\end{split}$ (3.49)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{2})|=&|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{1},0)|+|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|+|\widetilde{X}(z_{2},0)-\widetilde{X}(z_{2},t_{1})|\\ &+|\widetilde{X}(z_{2},t_{1})-\widetilde{X}(z_{2},t_{2})|\\ \leq&\sup_{\Omega}|\widetilde{X}_{t}|\cdot t_{1}+\frac{2}{\overline{K}}(\overline{M_{1}}+\sup_{\Omega}|\widetilde{W}|)|z_{2}-z_{1}|^{1/3}+\sup_{\Omega}|\widetilde{X}_{t}|\cdot t_{1}+\sup_{\Omega}|\widetilde{X}_{t}|\cdot|t_{1}-t_{2}|\\ \leq&\left(3\overline{M_{1}}+\frac{2}{\overline{K}}(\overline{M_{1}}+\sup_{\Omega}|\widetilde{W}|)\right)|(z_{1},t_{1})-(z_{2},t_{2})|^{1/3}.\end{split}$ (3.50)
@@
The inequalities above follow from Lemma 3.1 together with the explicit formulas for $f$, $g$ and $H_{ij}$ given in (3.24) and (3.9). Furthermore, by Lemma 3.3, there exists a uniform positive constant $M_{2}$, depending on the parameter $\nu$, such that
$|\widetilde{\Xi}|,|\widetilde{\Theta}|,t^{\nu}|\widetilde{X}_{z}|,t^{\nu}|\widetilde{Y}_{z}|\leq\overline{M_{2}}.$ (3.52)
@@
$\begin{split}|\widetilde{W}(z_{1},t_{m})-\widetilde{W}(z_{2},t_{m})|&\leq\frac{1}{t_{m}}\left(\left|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})\right|+\left|\widetilde{Y}(z_{1},t_{m})-\widetilde{Y}(z_{2},t_{m})\right|\right)\\ &\leq t_{m}^{-1-\nu}\left(\left|t_{m}^{\nu}\widetilde{X}(z^{\prime},t_{m})\right|\cdot|z_{1}-z_{2}|+\left|t_{m}^{\nu}\widetilde{Y}(z^{\prime\prime},t_{m})\right|\cdot|z_{1}-z_{2}|\right)\\ &\leq 2\overline{M_{2}}t_{m}^{-1-\nu}|z_{1}-z_{2}|\leq 2\overline{M_{2}}\;\overline{K}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}.\end{split}$ (3.56)
@@
$|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|,|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|,|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|\leq\overline{M_{3}}|z_{1}-z_{2}|^{1/3},$ (3.57)
for a certain uniform positive constant $\overline{M_{3}}$. Consequently, combining (3.51) and (3.57), we conclude that for $i=1,2$,
$\begin{split}\left|(H_{1i}-H_{2i})(z_{1},t)-(H_{1i}-H_{2i})(z_{2},t)\right|&\leq|H_{1i\widetilde{X}}-H_{2i\widetilde{X}}|\cdot|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|+|H_{1i\widetilde{Y}}-H_{2i\widetilde{Y}}|\\ &\cdot|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|+|H_{1ic}-H_{2ic}|\cdot c_{z}\cdot|z_{1}-z_{2}|\\ &\leq 4M^{*}\overline{M_{3}}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|.\end{split}$
@@
$\begin{split}|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|\leq&2\overline{M_{2}}\;\overline{K}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+\int_{0}^{t_{m}}\overline{M_{2}}(2M^{*}+1)t^{1-\nu}dt\\ &+\int_{0}^{t_{m}}2M^{*}\overline{M_{3}}|z_{1}-z_{2}|^{1/3}dt+\int_{0}^{t_{m}}\left(M^{*}+\frac{1}{2}\right)\left(4M^{*}\overline{M_{3}}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|\right)dt\\ \leq&2\overline{M_{2}}\;\overline{K}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+\frac{2M^{*}+1}{2-\nu}\overline{M_{2}}t_{m}^{2-\nu}+2M^{*}\overline{M_{3}}|z_{1}-z_{2}|^{1/3}t_{m}\\ &+(M^{*}+1)\left(4M^{*}\overline{M_{3}}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|\right)t_{m}\\ \leq&\overline{M_{4}}|z_{1}-z_{2}|^{\frac{2-\nu}{3}},\end{split}$ (3.58)
for a uniform positive constant $\overline{M_{4}}$ that depends on $\nu$. In view of (3.58), the same argument employed in Lemma 3.5 can be repeated to conclude that the function $\widetilde{W}$ is uniformly $C^{(2-\nu)/3}$ continuous throughout the entire domain $\Omega$.
@@
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|=&|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})|+\int_{0}^{t_{m}}|H_{12}(z_{1},t)-H_{12}(z_{2},t)|tdt\\ +\int_{0}^{t_{m}}&\left(\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{1},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{1},t)|+\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{2},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{2},t)|\right)t^{2-\nu}dt\\ +\int_{0}^{t_{m}}&\left(|2tH_{11}(z_{1},t)+1|\cdot|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|+2t|\widetilde{W}|+2t|\widetilde{W}|\cdot|H_{11}(z_{1},t)-H_{11}(z_{2},t)|\right)dt.\end{split}$ (3.60)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})|&=|\widetilde{X}_{z}(z^{\prime},t_{m})|\cdot|z_{1}-z_{2}|\\ &\leq\overline{M_{2}}t_{m}^{-\nu}|z_{1}-z_{2}|\leq\overline{M_{2}}\;\overline{K_{2}}^{\nu}|z_{1}-z_{2}|^{1-\nu}.\end{split}$ (3.61)
@@
$\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{1},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{1},t)|+\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{2},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{2},t)|\leq 2M^{*}\overline{M_{2}}.$
@@
$\begin{split}|H_{1i}(z_{1},t)-H_{1i}(z_{2},t)|\leq&|H_{1i\widetilde{X}}|\cdot|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|+|H_{1i\widetilde{Y}}|\cdot|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|\\ &+|H_{1i\tilde{c}}|\cdot|\tilde{c}(z_{1},t)-\tilde{c}(z_{2},t)|\\ \leq&\left(2\overline{M_{2}}t^{-\nu}+1\right)M^{*}|z_{1}-z_{2}|.\end{split}$
@@
$|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|\leq\overline{M_{5}}|z_{1}-z_{2}|^{\frac{2-\nu}{3}},$ (3.62)
for a uniform positive constant $\overline{M_{5}}$ depending on $\nu$. Employing (3.61)-(3.62) into (3.60) and then using (3.46) and (3.51), we conclude that
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|&\leq\overline{M_{2}}\;\overline{K}^{\nu}|z_{1}-z_{2}|^{1-\nu}+2M^{*}\overline{M_{2}}\int_{0}^{t_{m}}t^{2-\nu}dt\\ &+\int_{0}^{t_{m}}\left(M^{*}\overline{M_{5}}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+t(2M^{*}+1)(2\overline{M_{2}}t^{-\nu}+1)M^{*}|z_{1}-z_{2}|\right)dt\\ &=\overline{M_{2}}\;\overline{K}^{\nu}|z_{1}-z_{2}|^{1-\nu}+\frac{2M^{*}\overline{M_{2}}}{3-\nu}t_{m}^{3-\nu}\\ &+M^{*}\overline{M_{5}}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}t_{m}+(2M^{*}+1)(2\overline{M_{2}}+1)M^{*}|z_{1}-z_{2}|\leq\overline{M_{6}}|z_{1}-z_{2}|^{1-\nu},\end{split}$ (3.63)
for a uniform positive constant $\overline{M_{6}}$. Based on (3.63), arguments analogous to those used in Lemma 3.5 can be applied to conclude that the function $\widetilde{X}$ is uniformly $C^{1-\nu}$ continuous over the entire domain $\Omega$. By the same reasoning, the function $\widetilde{Y}$ also enjoys uniform $C^{1-\nu}$ continuity. This completes the proof of the lemma. ∎

### v2

@@ 3.5 Regularity of the solution
$K_{1}t_{m}\leq|z_{2}-z_{1}|^{1/3}\leq K_{2}t_{m},$ (3.46)
for some uniform positive constants $K_{1}$ and $K_{2}$. Since $\widetilde{W}$ is uniformly bounded, we may then apply (3.23) together with Lemma 3.3 to obtain
$|\tilde{\partial}_{+}\widetilde{X}|,|\tilde{\partial}_{-}\widetilde{Y}|,|\widetilde{X}_{t}|,|\widetilde{Y}_{t}|,t^{1/2}|\widetilde{X}_{z}|,t^{1/2}|\widetilde{Y}_{z}|\leq{a_{1}},$ (3.47)
for some positive constant $a_{1}>0$. Therefore, by Corollary 3.2 together with (3.46) and (3.47), we obtain
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|&=|\widetilde{Y}(z_{1},0)-\widetilde{X}(z_{2},0)|\\ &\leq|\widetilde{Y}(z_{1},0)-\widetilde{Y}(z_{m},t_{m})|+|\widetilde{Y}(z_{m},t_{m})-\widetilde{X}(z_{m},t_{m})|+|\widetilde{X}(z_{m},t_{m})-\widetilde{X}(z_{2},0)|\\ &\leq a_{1}t_{m}+2\sup_{\Omega}|\widetilde{W}|\cdot t_{m}+a_{1}t_{m}\\ &\leq\frac{2}{K_{2}}(a_{1}+\sup_{\Omega}|\widetilde{W}|)|z_{2}-z_{1}|^{1/3}.\end{split}$ (3.48)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{2})|&=|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{1})|+|\widetilde{X}(z_{2},t_{1})-\widetilde{X}(z_{2},t_{2})|\\ &\leq\sup_{\Omega}|\widetilde{X}_{z}|\cdot|z_{1}-z_{2}|+\sup_{\Omega}|\widetilde{X}_{t}|\cdot|t_{1}-t_{2}|\\ &\leq\frac{a_{1}}{t^{1/2}}|z_{1}-z_{2}|+a_{1}|t_{1}-t_{2}|\\ &\leq a_{1}|z_{1}-z_{2}|^{1/2}+a_{1}|t_{1}-t_{2}|\\ &\leq 2a_{1}|(z_{1},t_{1})-(z_{2},t_{2})|^{1/3},\end{split}$ (3.49)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{2},t_{2})|=&|\widetilde{X}(z_{1},t_{1})-\widetilde{X}(z_{1},0)|+|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|+|\widetilde{X}(z_{2},0)-\widetilde{X}(z_{2},t_{1})|\\ &+|\widetilde{X}(z_{2},t_{1})-\widetilde{X}(z_{2},t_{2})|\\ \leq&\sup_{\Omega}|\widetilde{X}_{t}|\cdot t_{1}+\frac{2}{K_{2}}(a_{1}+\sup_{\Omega}|\widetilde{W}|)|z_{2}-z_{1}|^{1/3}+\sup_{\Omega}|\widetilde{X}_{t}|\cdot t_{1}+\sup_{\Omega}|\widetilde{X}_{t}|\cdot|t_{1}-t_{2}|\\ \leq&\left(3a_{1}+\frac{2}{K_{2}}(a_{1}+\sup_{\Omega}|\widetilde{W}|)\right)|(z_{1},t_{1})-(z_{2},t_{2})|^{1/3}.\end{split}$ (3.50)
@@
The inequalities above follow from Lemma 3.1 together with the explicit formulas for $f$, $g$ and $H_{ij}$ given in (3.24) and (3.9). Furthermore, by Lemma 3.3, there exists a uniform positive constant $a_{2}$, depending on the parameter $\nu$, such that
$|\widetilde{\Xi}|,|\widetilde{\Theta}|,t^{\nu}|\widetilde{X}_{z}|,t^{\nu}|\widetilde{Y}_{z}|\leq a_{2}.$ (3.52)
@@
$\begin{split}|\widetilde{W}(z_{1},t_{m})-\widetilde{W}(z_{2},t_{m})|&\leq\frac{1}{t_{m}}\left(\left|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})\right|+\left|\widetilde{Y}(z_{1},t_{m})-\widetilde{Y}(z_{2},t_{m})\right|\right)\\ &\leq t_{m}^{-1-\nu}\left(\left|t_{m}^{\nu}\widetilde{X}(z^{\prime},t_{m})\right|\cdot|z_{1}-z_{2}|+\left|t_{m}^{\nu}\widetilde{Y}(z^{\prime\prime},t_{m})\right|\cdot|z_{1}-z_{2}|\right)\\ &\leq 2a_{2}t_{m}^{-1-\nu}|z_{1}-z_{2}|\leq 2a_{2}\;K_{2}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}.\end{split}$ (3.56)
@@
$|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|,|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|,|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|\leq a_{3}|z_{1}-z_{2}|^{1/3},$ (3.57)
for a certain uniform positive constant $a_{3}$. Consequently, combining (3.51) and (3.57), we conclude that for $i=1,2$,
$\begin{split}\left|(H_{1i}-H_{2i})(z_{1},t)-(H_{1i}-H_{2i})(z_{2},t)\right|&\leq|H_{1i\widetilde{X}}-H_{2i\widetilde{X}}|\cdot|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|+|H_{1i\widetilde{Y}}-H_{2i\widetilde{Y}}|\\ &\cdot|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|+|H_{1ic}-H_{2ic}|\cdot c_{z}\cdot|z_{1}-z_{2}|\\ &\leq 4M^{*}a_{3}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|.\end{split}$
@@
$\begin{split}|\widetilde{W}(z_{1},0)-\widetilde{W}(z_{2},0)|\leq&2a_{2}\;K_{2}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+\int_{0}^{t_{m}}a_{2}(2M^{*}+1)t^{1-\nu}dt\\ &+\int_{0}^{t_{m}}2M^{*}a_{3}|z_{1}-z_{2}|^{1/3}dt+\int_{0}^{t_{m}}\left(M^{*}+\frac{1}{2}\right)\left(4M^{*}a_{3}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|\right)dt\\ \leq&2a_{2}\;K_{2}^{1+\nu}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+\frac{2M^{*}+1}{2-\nu}a_{2}t_{m}^{2-\nu}+2M^{*}a_{3}|z_{1}-z_{2}|^{1/3}t_{m}\\ &+(M^{*}+1)\left(4M^{*}a_{3}|z_{1}-z_{2}|^{1/3}+2M^{*}|z_{1}-z_{2}|\right)t_{m}\\ \leq&a_{4}|z_{1}-z_{2}|^{\frac{2-\nu}{3}},\end{split}$ (3.58)
for a uniform positive constant $a_{4}$ that depends on $\nu$. In view of (3.58), the same argument employed in Lemma 3.5 can be repeated to conclude that the function $\widetilde{W}$ is uniformly $C^{(2-\nu)/3}$ continuous throughout the entire domain $\Omega$.
@@
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|\leq&|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})|+\int_{0}^{t_{m}}|H_{12}(z_{1},t)-H_{12}(z_{2},t)|tdt\\ +\int_{0}^{t_{m}}&\left(\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{1},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{1},t)|+\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{2},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{2},t)|\right)t^{2-\nu}dt\\ +\int_{0}^{t_{m}}&\left(|2tH_{11}(z_{1},t)+1|\cdot|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|+2t|\widetilde{W}|+2t|\widetilde{W}|\cdot|H_{11}(z_{1},t)-H_{11}(z_{2},t)|\right)dt.\end{split}$ (3.60)
@@
$\begin{split}|\widetilde{X}(z_{1},t_{m})-\widetilde{X}(z_{2},t_{m})|&=|\widetilde{X}_{z}(z^{\prime},t_{m})|\cdot|z_{1}-z_{2}|\leq a_{2}t_{m}^{-\nu}|z_{1}-z_{2}|\leq a_{2}{K_{2}}^{\nu}|z_{1}-z_{2}|^{1-\nu}.\end{split}$ (3.61)
@@
$\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{1},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{1},t)|+\left|\frac{cf\widetilde{Y}}{1+tg\widetilde{Y}}(z_{2},t)\right|\cdot|t^{\nu}\widetilde{X}_{z}(z_{2},t)|\leq 2M^{*}a_{2}.$
@@
$\begin{split}|H_{1i}(z_{1},t)-H_{1i}(z_{2},t)|\leq&|H_{1i\widetilde{X}}|\cdot|\widetilde{X}(z_{1},t)-\widetilde{X}(z_{2},t)|+|H_{1i\widetilde{Y}}|\cdot|\widetilde{Y}(z_{1},t)-\widetilde{Y}(z_{2},t)|\\ &+|H_{1i\tilde{c}}|\cdot|\tilde{c}(z_{1},t)-\tilde{c}(z_{2},t)|\\ \leq&\left(2a_{2}t^{-\nu}+1\right)M^{*}|z_{1}-z_{2}|.\end{split}$
@@
$|\widetilde{W}(z_{1},t)-\widetilde{W}(z_{2},t)|\leq a_{5}|z_{1}-z_{2}|^{\frac{2-\nu}{3}},$ (3.62)
for a uniform positive constant $a_{5}$ depending on $\nu$. Employing (3.61)-(3.62) into (3.60) and then using (3.46) and (3.51), we conclude that
$\begin{split}|\widetilde{X}(z_{1},0)-\widetilde{X}(z_{2},0)|&\leq a_{2}K_{2}^{\nu}|z_{1}-z_{2}|^{1-\nu}+2M^{*}a_{2}\int_{0}^{t_{m}}t^{2-\nu}dt\\ &+M^{*}\int_{0}^{t_{m}}\left(a_{5}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}+t(2M^{*}+1)(2a_{2}t^{-\nu}+1)|z_{1}-z_{2}|\right)dt\\ &=a_{2}K_{2}^{\nu}|z_{1}-z_{2}|^{1-\nu}+\frac{2M^{*}a_{2}}{3-\nu}t_{m}^{3-\nu}+M^{*}a_{5}|z_{1}-z_{2}|^{\frac{2-\nu}{3}}t_{m}+(2M^{*}+1)(2a_{2}+1)M^{*}|z_{1}-z_{2}|\\ &\leq a_{6}|z_{1}-z_{2}|^{1-\nu},\end{split}$ (3.63)
for a uniform positive constant $a_{6}$. Based on (3.63), arguments analogous to those used in Lemma 3.5 can be applied to conclude that the function $\widetilde{X}$ is uniformly $C^{1-\nu}$ continuous over the entire domain $\Omega$. By the same reasoning, the function $\widetilde{Y}$ also enjoys uniform $C^{1-\nu}$ continuity. This completes the proof of the lemma. ∎